OpenStudy (anonymous):

Write an equation for the line that passes through (0, -2) and whose slope is ½. y = 1/2x + 2 y = -1/2x -2 y = 1/2x – 2 y = -2x + ½

OpenStudy (whpalmer4):

Use the formula \(y - y_0 = m(x-x_0)\) to write the equation of a line passing through the point \(x_0,y_0\) with slope \(m\). Then solve for y and match with your answer choices.

OpenStudy (anonymous):

Oh okay I get it

OpenStudy (texaschic101):

y - y1 = m(x - x1) set 1(0,-2) slope 1/2 Let me get you started... sub in your info into the formula.. y - (-2) = 1/2(x - 0) (distribute through the parenthesis) y + 2 = 1/2x Now, put this into the y = mx + b formula and you will have your answer

OpenStudy (anonymous):

:) thanks

OpenStudy (jdoe0001):

http://www.teachhandson.com/calculationb.gif